[Differential equation]

Solve the following ODE by any metod:

? 2? ′′ + 3?? ′ − 7? = 0
Answer #1

Given differential equation is ry + 3xy-7y=040 Here Y is differentiated with respec t to x. dy + 3x dy - 79 =O © (By usingecon How equation ® becomes Dat9 +304-79=0 02_D) y + 3194– 7y=0 (028 +39 -7) y=0 Co+20-774703 Now Euler-Cauchys equation RedThe Auriliary equation is fo=0 = 3²+29_7=0 To find the roots by - using Quadratic formula D = -b Iſ6²-4ac 2a Here La=i, b=2,cDL-27 14x8 2 ♡ = -2 + 218 D = -1 + 18 D = 7+38 02-1-VE zt +14x2, D = 7-14 82 34/7272,D=-1-202 : The roots are D = 1+2V2, -1-2& c = 4, & 1+2 V2Jt tte C+2V2) tuo - C+ 2 VEJE and Yp = particular integralis p. I FED) = .con 042D-7 :lypropy1. Hence the general soluation is y = 4 typ =c, € +2.00€ tse (1+202) sei +2023€ Cla+212)t - the replace it by loge =c, (+2V2)tex Yze, 4-1+207) (C+2V734 Yz4, 2614203) 72 (Stavi] y= (-142V2) ę è CY (1+212) Hence 202) 19 = cf1212) C 2 (442V2) Which is r